OpenStudy (anonymous):

What is the resulting equation when the expression for y in the second equation is substituted into the first equation? 3x + y = 1 y = 6 - 4x

OpenStudy (anonymous):

3x + (6-4x) = 1 3x + 6 - 4x = 1 6 - x = 1 -x = 1-6 -x = -5 x = 5 y=6-4x y=6-4(5) y=6-20 y=-14

OpenStudy (texaschic101):

the equation is 3x + (6 - 4x) = 1